# Separable Mixture of Low-Rank Adaptation for Continual Visual Instruction Tuning

### v1

@@ 6 Conclusion
In this paper, we identify a dual form of catastrophic forgetting in CVIT and propose an efficient SMoLoRA method. SMoLoRA leverages separable routing to dynamically assign appropriate LoRA blocks to different tasks, thereby preserving the model’s visual understanding and instruction following capabilities, mitigating the dual catastrophic forgetting. Additionally, we introduce a new CVIT benchmark that evaluates not only model accuracy and forgetting on standard tasks but also its ability to generalize to unseen tasks and adapt to diverse instructions. Extensive experiments on the CVIT benchmark demonstrate the superior performance of our SMoLoRA over existing methods.
Appendix

### v2

@@ 6 Conclusion
In this paper, we identify a dual form of catastrophic forgetting in CVIT and propose an effective SMoLoRA method. SMoLoRA leverages separable routing to dynamically assign appropriate LoRA blocks to different tasks, thereby preserving the model’s visual understanding and instruction following capabilities, mitigating the dual catastrophic forgetting. Additionally, we introduce a new CVIT benchmark that evaluates not only model accuracy and forgetting on standard tasks but also its ability to generalize to unseen tasks and adapt to diverse instructions. Extensive experiments on the CVIT benchmark demonstrate the superior performance of our SMoLoRA over existing methods.
Supplementary Material
@@ D. Details of CVIT Benchmark
E. Additional Case Studies
As shown in Fig. 9, we present additional cases to validate the efficacy of our proposed SMoLoRA. Individual modules (VU and IF) are only capable of resolving their respective forgetting problems, whereas the combination of both can simultaneously mitigate the dual catastrophic forgetting.
@@ E. Additional Case Studies
Figure 9: Additional Case Studies on the Effectiveness of Separation in SMoLoRA.